\begin{thebibliography}{99}

%\cite{Klinkhamer:1984di}
\bibitem{Klinkhamer:1984di} 
  F.~R.~Klinkhamer and N.~S.~Manton,
  %``A Saddle Point Solution in the Weinberg-Salam Theory,''
  Phys.\ Rev.\ D {\bf 30}, 2212 (1984).
  %doi:10.1103/PhysRevD.30.2212
  %%CITATION = doi:10.1103/PhysRevD.30.2212;%%
  %1173 citations counted in INSPIRE as of 13 Dec 2016
  
%\cite{Dashen:1974ck}
\bibitem{Dashen:1974ck} 
  R.~F.~Dashen, B.~Hasslacher and A.~Neveu,
  %``Nonperturbative Methods and Extended Hadron Models in Field Theory. 3. Four-Dimensional Nonabelian Models,''
  Phys.\ Rev.\ D {\bf 10}, 4138 (1974).
  %doi:10.1103/PhysRevD.10.4138
  %%CITATION = doi:10.1103/PhysRevD.10.4138;%%
  %328 citations counted in INSPIRE as of 25 Jan 2017
  
%\cite{Soni:1980ps}
\bibitem{Soni:1980ps} 
  V.~Soni,
  %``Possible Classical Solutions In The Weinberg-salam Model,''
  Phys.\ Lett.\  {\bf 93B}, 101 (1980).
  %doi:10.1016/0370-2693(80)90104-5
  %%CITATION = doi:10.1016/0370-2693(80)90104-5;%%
  %37 citations counted in INSPIRE as of 25 Jan 2017
  
%\cite{Boguta:1983xs}
\bibitem{Boguta:1983xs} 
  J.~Boguta,
  %``Can Nuclear Interactions Be Long Ranged?,''
  Phys.\ Rev.\ Lett.\  {\bf 50}, 148 (1983).
 % doi:10.1103/PhysRevLett.50.148
  %%CITATION = doi:10.1103/PhysRevLett.50.148;%%
  %82 citations counted in INSPIRE as of 25 Jan 2017
  
%\cite{Forgacs:1983yu}
\bibitem{Forgacs:1983yu} 
  P.~Forgacs and Z.~Horvath,
  %``Topology and Saddle Points in Field Theories,''
  Phys.\ Lett.\  {\bf 138B}, 397 (1984).
  %%doi:10.1016/0370-2693(84)91926-9
  %%CITATION = %doi:10.1016/0370-2693(84)91926-9;%%
  %52 citations counted in INSPIRE as of 25 Jan 2017



%\cite{Sakharov:1967dj}
\bibitem{Sakharov:1967dj} 
  A.~D.~Sakharov,
  %``Violation of CP Invariance, c Asymmetry, and Baryon Asymmetry of the Universe,''
  Pisma Zh.\ Eksp.\ Teor.\ Fiz.\  {\bf 5}, 32 (1967)
  [JETP Lett.\  {\bf 5}, 24 (1967)]
  [Sov.\ Phys.\ Usp.\  {\bf 34}, 392 (1991)]
  [Usp.\ Fiz.\ Nauk {\bf 161}, 61 (1991)].
  %doi:10.1070/PU1991v034n05ABEH002497
  %%CITATION = %doi:10.1070/PU1991v034n05ABEH002497;%%
  %2523 citations counted in INSPIRE as of 13 Dec 2016

%\cite{Riotto:1999yt}
\bibitem{Riotto:1999yt}
  A.~Riotto and M.~Trodden,
  %``Recent progress in baryogenesis,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 49} (1999) 35
  %doi:10.1146/annurev.nucl.49.1.35
  [hep-ph/9901362].
  %%CITATION = %doi:10.1146/annurev.nucl.49.1.35;%%
  %459 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Cohen:1993nk}
\bibitem{Cohen:1993nk} 
  A.~G.~Cohen, D.~B.~Kaplan and A.~E.~Nelson,
  %``Progress in electroweak baryogenesis,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 43}, 27 (1993)
  %doi:10.1146/annurev.ns.43.120193.000331
  [hep-ph/9302210].
  %%CITATION = %doi:10.1146/annurev.ns.43.120193.000331;%%
  %744 citations counted in INSPIRE as of 06 Dec 2016
 
%\cite{Rubakov:1996vz}
\bibitem{Rubakov:1996vz} 
  V.~A.~Rubakov and M.~E.~Shaposhnikov,
  %``Electroweak baryon number nonconservation in the early universe and in high-energy collisions,''
  Usp.\ Fiz.\ Nauk {\bf 166}, 493 (1996)
  [Phys.\ Usp.\  {\bf 39}, 461 (1996)]
  %doi:10.1070/PU1996v039n05ABEH000145
  [hep-ph/9603208].
  %%CITATION = %doi:10.1070/PU1996v039n05ABEH000145;%%
  %
  
  
%\cite{Kharzeev:2007jp}
\bibitem{Kharzeev:2007jp} 
  D.~E.~Kharzeev, L.~D.~McLerran and H.~J.~Warringa,
  %``The Effects of topological charge change in heavy ion collisions: 'Event by event P and CP violation',''
  Nucl.\ Phys.\ A {\bf 803}, 227 (2008)
  %doi:10.1016/j.nuclphysa.2008.02.298
  [arXiv:0711.0950 [hep-ph]].
  %%CITATION = %doi:10.1016/j.nuclphysa.2008.02.298;%%
  %822 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Fukushima:2008xe}
\bibitem{Fukushima:2008xe} 
  K.~Fukushima, D.~E.~Kharzeev and H.~J.~Warringa,
  %``The Chiral Magnetic Effect,''
  Phys.\ Rev.\ D {\bf 78}, 074033 (2008)
  %doi:10.1103/PhysRevD.78.074033
  [arXiv:0808.3382 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevD.78.074033;%%
  %755 citations counted in INSPIRE as of 06 Dec 2016

%\cite{Kharzeev:2015znc}
\bibitem{Kharzeev:2015znc} 
  D.~E.~Kharzeev, J.~Liao, S.~A.~Voloshin and G.~Wang,
  %``Chiral magnetic and vortical effects in high-energy nuclear collisions—A status report,''
  Prog.\ Part.\ Nucl.\ Phys.\  {\bf 88}, 1 (2016)
  %doi:10.1016/j.ppnp.2016.01.001
  [arXiv:1511.04050 [hep-ph]].
  %%CITATION = %doi:10.1016/j.ppnp.2016.01.001;%%
  %70 citations counted in INSPIRE as of 06 Dec 2016
  
  
%\cite{Skokov:2016yrj}
\bibitem{Skokov:2016yrj} 
  V.~Skokov, P.~Sorensen, V.~Koch, S.~Schlichting, J.~Thomas, S.~Voloshin, G.~Wang and H.~U.~Yee,
  %``Chiral Magnetic Effect Task Force Report,''
  arXiv:1608.00982 [nucl-th].
  %%CITATION = ARXIV:1608.00982;%%
  %9 citations counted in INSPIRE as of 06 Dec 2016
  
  
  
%\cite{Skokov:2009qp}
\bibitem{Skokov:2009qp} 
  V.~Skokov, A.~Y.~Illarionov and V.~Toneev,
  %``Estimate of the magnetic field strength in heavy-ion collisions,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 24}, 5925 (2009)
  %doi:10.1142/S0217751X09047570
  [arXiv:0907.1396 [nucl-th]].
  %%CITATION = %doi:10.1142/S0217751X09047570;%%
  %415 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Deng:2012pc}
\bibitem{Deng:2012pc} 
  W.~T.~Deng and X.~G.~Huang,
  %``Event-by-event generation of electromagnetic fields in heavy-ion collisions,''
  Phys.\ Rev.\ C {\bf 85}, 044907 (2012)
  %doi:10.1103/PhysRevC.85.044907
  [arXiv:1201.5108 [nucl-th]].
  %%CITATION = %doi:10.1103/PhysRevC.85.044907;%%
  %186 citations counted in INSPIRE as of 06 Dec 2016
  
  
%\cite{Lappi:2006fp}
\bibitem{Lappi:2006fp} 
  T.~Lappi and L.~McLerran,
  %``Some features of the glasma,''
  Nucl.\ Phys.\ A {\bf 772}, 200 (2006)
  doi:10.1016/j.nuclphysa.2006.04.001
  [hep-ph/0602189].
  %%CITATION = doi:10.1016/j.nuclphysa.2006.04.001;%%
  %386 citations counted in INSPIRE as of 11 Dec 2017
  
  %\cite{Gelis:2006dv}
\bibitem{Gelis:2006dv} 
  F.~Gelis and R.~Venugopalan,
  %``Three lectures on multi-particle production in the glasma,''
  Acta Phys.\ Polon.\ B {\bf 37}, 3253 (2006)
  [hep-ph/0611157].
  %%CITATION = HEP-PH/0611157;%%
  %49 citations counted in INSPIRE as of 11 Dec 2017
  
  %\cite{Blaizot:2016qgz}
\bibitem{Blaizot:2016qgz} 
  J.~P.~Blaizot,
  %``High gluon densities in heavy ion collisions,''
  Rept.\ Prog.\ Phys.\  {\bf 80}, no. 3, 032301 (2017)
  doi:10.1088/1361-6633/aa5435
  [arXiv:1607.04448 [hep-ph]].
  %%CITATION = doi:10.1088/1361-6633/aa5435;%%
  %8 citations counted in INSPIRE as of 11 Dec 2017



  %\cite{McLerran:1993ni}
\bibitem{McLerran:1993ni} 
  L.~D.~McLerran and R.~Venugopalan,
  %``Computing quark and gluon distribution functions for very large nuclei,''
  Phys.\ Rev.\ D {\bf 49}, 2233 (1994)
  %doi:10.1103/PhysRevD.49.2233
  [hep-ph/9309289].
  %%CITATION = %doi:10.1103/PhysRevD.49.2233;%%
  %1699 citations counted in INSPIRE as of 05 Dec 2017
  
  %\cite{McLerran:1993ka}
\bibitem{McLerran:1993ka} 
  L.~D.~McLerran and R.~Venugopalan,
  %``Gluon distribution functions for very large nuclei at small transverse momentum,''
  Phys.\ Rev.\ D {\bf 49}, 3352 (1994)
  %doi:10.1103/PhysRevD.49.3352
  [hep-ph/9311205].
  %%CITATION = %doi:10.1103/PhysRevD.49.3352;%%
  %1252 citations counted in INSPIRE as of 05 Dec 2017
  
  %\cite{Gelis:2010nm}
\bibitem{Gelis:2010nm} 
  F.~Gelis, E.~Iancu, J.~Jalilian-Marian and R.~Venugopalan,
  %``The Color Glass Condensate,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 60}, 463 (2010)
  %doi:10.1146/annurev.nucl.010909.083629
  [arXiv:1002.0333 [hep-ph]].
  %%CITATION = %doi:10.1146/annurev.nucl.010909.083629;%%
  %616 citations counted in INSPIRE as of 05 Dec 2017  
  
  

%\cite{Berges:2013eia}
\bibitem{Berges:2013eia} 
  J.~Berges, K.~Boguslavski, S.~Schlichting and R.~Venugopalan,
  %``Turbulent thermalization process in heavy-ion collisions at ultrarelativistic energies,''
  Phys.\ Rev.\ D {\bf 89}, no. 7, 074011 (2014)
  doi:10.1103/PhysRevD.89.074011
  [arXiv:1303.5650 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.89.074011;%%
  %134 citations counted in INSPIRE as of 11 Dec 2017
  
  %\cite{Berges:2013fga}
\bibitem{Berges:2013fga} 
  J.~Berges, K.~Boguslavski, S.~Schlichting and R.~Venugopalan,
  %``Universal attractor in a highly occupied non-Abelian plasma,''
  Phys.\ Rev.\ D {\bf 89}, no. 11, 114007 (2014)
  doi:10.1103/PhysRevD.89.114007
  [arXiv:1311.3005 [hep-ph]].
  %%CITATION = doi:10.1103/PhysRevD.89.114007;%%
  %103 citations counted in INSPIRE as of 11 Dec 2017
  
  %\cite{Gelis:2016upa}
\bibitem{Gelis:2016upa} 
  F.~Gelis and B.~Schenke,
  %``Initial State Quantum Fluctuations in the Little Bang,''
  Ann.\ Rev.\ Nucl.\ Part.\ Sci.\  {\bf 66}, 73 (2016)
  doi:10.1146/annurev-nucl-102115-044651
  [arXiv:1604.00335 [hep-ph]].
  %%CITATION = doi:10.1146/annurev-nucl-102115-044651;%%
  %10 citations counted in INSPIRE as of 11 Dec 2017  
  %\cite{Berges:2012ev}
  
    
%\cite{Mace:2016svc}
\bibitem{Mace:2016svc} 
  M.~Mace, S.~Schlichting and R.~Venugopalan,
  %``Off-equilibrium sphaleron transitions in the Glasma,''
  Phys.\ Rev.\ D {\bf 93}, no. 7, 074036 (2016)
  %doi:10.1103/PhysRevD.93.074036
  [arXiv:1601.07342 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevD.93.074036;%%
  %9 citations counted in INSPIRE as of 06 Dec 2016 
  
%\cite{Moore:2010jd}
\bibitem{Moore:2010jd} 
  G.~D.~Moore and M.~Tassler,
  %``The Sphaleron Rate in SU(N) Gauge Theory,''
  JHEP {\bf 1102}, 105 (2011)
  %doi:10.1007/JHEP02(2011)105
  [arXiv:1011.1167 [hep-ph]].
  %%CITATION = %doi:10.1007/JHEP02(2011)105;%%
  %49 citations counted in INSPIRE as of 06 Dec 2016
  
  
%\cite{Mueller:2016ven}
\bibitem{Mueller:2016ven} 
  N.~Mueller, S.~Schlichting and S.~Sharma,
  %``Chiral magnetic effect and anomalous transport from real-time lattice simulations,''
  Phys.\ Rev.\ Lett.\  {\bf 117}, no. 14, 142301 (2016)
  %doi:10.1103/PhysRevLett.117.142301
  [arXiv:1606.00342 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevLett.117.142301;%%
  %5 citations counted in INSPIRE as of 09 Dec 2016
  
  %\cite{Mace:2016shq}
\bibitem{Mace:2016shq} 
  M.~Mace, N.~Mueller, S.~Schlichting and S.~Sharma,
  %``Non-equilibrium study of the Chiral Magnetic Effect from real-time simulations with dynamical fermions,''
  arXiv:1612.02477 [hep-lat].
  %%CITATION = ARXIV:1612.02477;%%
  
  %\cite{Mace:2017wcl}
\bibitem{Mace:2017wcl} 
  M.~Mace, N.~Mueller, S.~Schlichting and S.~Sharma,
  %``Simulating chiral magnetic effect and anomalous transport phenomena in the pre-equilibrium stages of heavy-ion collisions,''
  Nucl.\ Phys.\ A {\bf 967}, 752 (2017)
  %doi:10.1016/j.nuclphysa.2017.05.040
  [arXiv:1704.05887 [hep-lat]].
  %%CITATION = %doi:10.1016/j.nuclphysa.2017.05.040;%%
  
  
  
%\cite{Son:2009tf}
\bibitem{Son:2009tf} 
  D.~T.~Son and P.~Surowka,
  %``Hydrodynamics with Triangle Anomalies,''
  Phys.\ Rev.\ Lett.\  {\bf 103}, 191601 (2009)
  %doi:10.1103/PhysRevLett.103.191601
  [arXiv:0906.5044 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevLett.103.191601;%%
  %388 citations counted in INSPIRE as of 13 Dec 2016
  
  
%\cite{Nair:2011mk}
\bibitem{Nair:2011mk} 
  V.~P.~Nair, R.~Ray and S.~Roy,
  %``Fluids, Anomalies and the Chiral Magnetic Effect: A Group-Theoretic Formulation,''
  Phys.\ Rev.\ D {\bf 86}, 025012 (2012)
  %doi:10.1103/PhysRevD.86.025012
  [arXiv:1112.4022 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.86.025012;%%
  %52 citations counted in INSPIRE as of 05 May 2017

  
   %\cite{Gursoy:2014aka}
\bibitem{Gursoy:2014aka} 
  U.~Gursoy, D.~Kharzeev and K.~Rajagopal,
  %``Magnetohydrodynamics, charged currents and directed flow in heavy ion collisions,''
  Phys.\ Rev.\ C {\bf 89}, no. 5, 054905 (2014)
  %doi:10.1103/PhysRevC.89.054905
  [arXiv:1401.3805 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevC.89.054905;%%
  %64 citations counted in INSPIRE as of 12 Dec 2016
  
%\cite{Hongo:2013cqa}
\bibitem{Hongo:2013cqa} 
  M.~Hongo, Y.~Hirono and T.~Hirano,
  %``First Numerical Simulations of Anomalous Hydrodynamics,''
  arXiv:1309.2823 [nucl-th].
  %%CITATION = ARXIV:1309.2823;%%
  %32 citations counted in INSPIRE as of 25 Jan 2017


  %\cite{Karabali:2014vla}
\bibitem{Karabali:2014vla} 
  D.~Karabali and V.~P.~Nair,
  %``Relativistic Particle and Relativistic Fluids: Magnetic Moment and Spin-Orbit Interactions,''
  Phys.\ Rev.\ D {\bf 90}, no. 10, 105018 (2014)
  %doi:10.1103/PhysRevD.90.105018
  [arXiv:1406.1551 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.90.105018;%%
  %11 citations counted in INSPIRE as of 20 Apr 2017

%\cite{Hirono:2014oda}
\bibitem{Hirono:2014oda} 
  Y.~Hirono, T.~Hirano and D.~E.~Kharzeev,
  %``The chiral magnetic effect in heavy-ion collisions from event-by-event anomalous hydrodynamics,''
  arXiv:1412.0311 [hep-ph].
  %%CITATION = ARXIV:1412.0311;%%
  %28 citations counted in INSPIRE as of 13 Dec 2016
  
%\cite{Yin:2015fca}
\bibitem{Yin:2015fca} 
  Y.~Yin and J.~Liao,
  %``Hydrodynamics with chiral anomaly and charge separation in relativistic heavy ion collisions,''
  Phys.\ Lett.\ B {\bf 756}, 42 (2016)
  %doi:10.1016/j.physletb.2016.02.065
  [arXiv:1504.06906 [nucl-th]].
  %%CITATION = %doi:10.1016/j.physletb.2016.02.065;%%
  %12 citations counted in INSPIRE as of 13 Dec 2016
  

  

%\cite{Berry:1984jv}
\bibitem{Berry:1984jv} 
  M.~V.~Berry,
  %``Quantal phase factors accompanying adiabatic changes,''
  Proc.\ Roy.\ Soc.\ Lond.\ A {\bf 392}, 45 (1984).
  %doi:10.1098/rspa.1984.0023
  %%CITATION = %doi:10.1098/rspa.1984.0023;%%
  %993 citations counted in INSPIRE as of 25 Jan 2017
  

%\cite{Son:2012wh}
\bibitem{Son:2012wh} 
  D.~T.~Son and N.~Yamamoto,
  %``Berry Curvature, Triangle Anomalies, and the Chiral Magnetic Effect in Fermi Liquids,''
  Phys.\ Rev.\ Lett.\  {\bf 109}, 181602 (2012)
  %doi:10.1103/PhysRevLett.109.181602
  [arXiv:1203.2697 [cond-mat.mes-hall]].
  %%CITATION = %doi:10.1103/PhysRevLett.109.181602;%%
  %130 citations counted in INSPIRE as of 11 Aug 2016
  
%\cite{Stephanov:2012ki}
\bibitem{Stephanov:2012ki} 
  M.~A.~Stephanov and Y.~Yin,
  %``Chiral Kinetic Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 109}, 162001 (2012)
  %doi:10.1103/PhysRevLett.109.162001
  [arXiv:1207.0747 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevLett.109.162001;%%
  %106 citations counted in INSPIRE as of 11 Aug 2016
  
%\cite{Son:2012zy}
\bibitem{Son:2012zy} 
  D.~T.~Son and N.~Yamamoto,
  %``Kinetic theory with Berry curvature from quantum field theories,''
  Phys.\ Rev.\ D {\bf 87}, no. 8, 085016 (2013)
  %doi:10.1103/PhysRevD.87.085016
  [arXiv:1210.8158 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.87.085016;%%
  %84 citations counted in INSPIRE as of 11 Aug 2016
 %\cite{Chen:2014cla}
 
  %\cite{Chen:2013iga}
\bibitem{Chen:2013iga}
  J.~W.~Chen, J.~y.~Pang, S.~Pu and Q.~Wang,
  %``Kinetic equations for massive Dirac fermions in electromagnetic field with non-Abelian Berry phase,''
  Phys.\ Rev.\ D {\bf 89}, no. 9, 094003 (2014)
  %doi:10.1103/PhysRevD.89.094003
  [arXiv:1312.2032 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.89.094003;%%
  %19 citations counted in INSPIRE as of 06 Jun 2016
 
\bibitem{Chen:2014cla} 
  J.~Y.~Chen, D.~T.~Son, M.~A.~Stephanov, H.~U.~Yee and Y.~Yin,
  %``Lorentz Invariance in Chiral Kinetic Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 113}, no. 18, 182302 (2014)
  %doi:10.1103/PhysRevLett.113.182302
  [arXiv:1404.5963 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevLett.113.182302;%%
  %42 citations counted in INSPIRE as of 11 Aug 2016 
  
  
  \bibitem{Stone:2013sga} 
  M.~Stone and V.~Dwivedi,
  %``Classical version of the non-Abelian gauge anomaly,''
  Phys.\ Rev.\ D {\bf 88}, no. 4, 045012 (2013)
  %doi:10.1103/PhysRevD.88.045012
  [arXiv:1305.1955 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.88.045012;%%
  %18 citations counted in INSPIRE as of 13 Aug 2016
  
 \bibitem{Dwivedi:2013dea} 
  V.~Dwivedi and M.~Stone,
  %``Classical chiral kinetic theory and anomalies in even space-time dimensions,''
  J.\ Phys.\ A {\bf 47}, 025401 (2013)
  %doi:10.1088/1751-8113/47/2/025401
  [arXiv:1308.4576 [hep-th]].
  %%CITATION = %doi:10.1088/1751-8113/47/2/025401;%%
  %12 citations counted in INSPIRE as of 13 Aug 2016
  
  
   
%\cite{Manuel:2014dza}
\bibitem{Manuel:2014dza} 
  C.~Manuel and J.~M.~Torres-Rincon,
  %``Chiral transport equation from the quantum Dirac Hamiltonian and the on-shell effective field theory,''
  Phys.\ Rev.\ D {\bf 90}, no. 7, 076007 (2014)
  %doi:10.1103/PhysRevD.90.076007
  [arXiv:1404.6409 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevD.90.076007;%%
  %28 citations counted in INSPIRE as of 19 Jan 2017

%\cite{Stone:2014fja}
\bibitem{Stone:2014fja} 
  M.~Stone, V.~Dwivedi and T.~Zhou,
  %``Berry Phase, Lorentz Covariance, and Anomalous Velocity for Dirac and Weyl Particles,''
  Phys.\ Rev.\ D {\bf 91}, no. 2, 025004 (2015)
  %doi:10.1103/PhysRevD.91.025004
  [arXiv:1406.0354 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevD.91.025004;%%
  %15 citations counted in INSPIRE as of 01 Oct 2016

  %\cite{Manuel:2015zpa}
\bibitem{Manuel:2015zpa} 
  C.~Manuel and J.~M.~Torres-Rincon,
  %``Dynamical evolution of the chiral magnetic effect: Applications to the quark-gluon plasma,''
  Phys.\ Rev.\ D {\bf 92}, no. 7, 074018 (2015)
  %doi:10.1103/PhysRevD.92.074018
  [arXiv:1501.07608 [hep-ph]].
  %%CITATION = %doi:10.1103/PhysRevD.92.074018;%%
  %24 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Chen:2015gta}
\bibitem{Chen:2015gta} 
  J.~Y.~Chen, D.~T.~Son and M.~A.~Stephanov,
  %``Collisions in Chiral Kinetic Theory,''
  Phys.\ Rev.\ Lett.\  {\bf 115}, no. 2, 021601 (2015)
  %doi:10.1103/PhysRevLett.115.021601
  [arXiv:1502.06966 [hep-th]].
  %%CITATION = %doi:10.1103/PhysRevLett.115.021601;%%
  %13 citations counted in INSPIRE as of 22 Nov 2016

\bibitem{Sun:2016nig} 
  Y.~Sun, C.~M.~Ko and F.~Li,
  %``Anomalous transport model study of chiral magnetic effects in heavy ion collisions,''
  Phys.\ Rev.\ C {\bf 94}, no. 4, 045204 (2016)
  %doi:10.1103/PhysRevC.94.045204
  [arXiv:1606.05627 [nucl-th]].
  %%CITATION = %doi:10.1103/PhysRevC.94.045204;%%

%\cite{Hidaka:2016yjf}
\bibitem{Hidaka:2016yjf} 
  Y.~Hidaka, S.~Pu and D.~L.~Yang,
  %``Relativistic Chiral Kinetic Theory from Quantum Field Theories,''
  arXiv:1612.04630 [hep-th].
  %%CITATION = ARXIV:1612.04630;%%
  
  
%\cite{Bodeker:1999ey,Litim:1999id}Akamatsu:2013pjd
\bibitem{Bodeker:1999ey} 
  D.~B\"{o}deker,
  %``From hard thermal loops to Langevin dynamics,''
  Nucl.\ Phys.\ B {\bf 559}, 502 (1999).

  \bibitem{Litim:1999id} 
  D.~F.~Litim and C.~Manuel,
  %``Effective transport equations for nonAbelian plasmas,''
  Nucl.\ Phys.\ B {\bf 562}, 237 (1999).


\bibitem{Akamatsu:2013pjd} 
  Y.~Akamatsu and N.~Yamamoto,
  %``Chiral Plasma Instabilities,''
  Phys.\ Rev.\ Lett.\  {\bf 111}, 052002 (2013).
  
  
  %\cite{Akamatsu:2014yza,Yamamoto:2015gzz,Akamatsu:2015kau}
\bibitem{Akamatsu:2014yza} 
  Y.~Akamatsu and N.~Yamamoto,
  %``Chiral Langevin theory for non-Abelian plasmas,''
  Phys.\ Rev.\ D {\bf 90}, no. 12, 125031 (2014)
  doi:10.1103/PhysRevD.90.125031
  [arXiv:1402.4174 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.90.125031;%%
  %25 citations counted in INSPIRE as of 11 Dec 2017

%\cite{Yamamoto:2015gzz,Akamatsu:2015kau}
\bibitem{Yamamoto:2015gzz} 
  N.~Yamamoto,
  %``Chiral transport of neutrinos in supernovae: Neutrino-induced fluid helicity and helical plasma instability,''
  Phys.\ Rev.\ D {\bf 93}, no. 6, 065017 (2016)
  doi:10.1103/PhysRevD.93.065017
  [arXiv:1511.00933 [astro-ph.HE]].
  %%CITATION = doi:10.1103/PhysRevD.93.065017;%%
  %29 citations counted in INSPIRE as of 11 Dec 2017
  
  %\cite{Akamatsu:2015kau}
\bibitem{Akamatsu:2015kau} 
  Y.~Akamatsu, A.~Rothkopf and N.~Yamamoto,
  %``Non-Abelian chiral instabilities at high temperature on the lattice,''
  JHEP {\bf 1603}, 210 (2016)
  doi:10.1007/JHEP03(2016)210
  [arXiv:1512.02374 [hep-ph]].
  %%CITATION = doi:10.1007/JHEP03(2016)210;%%
  %8 citations counted in INSPIRE as of 11 Dec 2017

%\cite{Yamamoto:2017uul}
\bibitem{Yamamoto:2017uul} 
  N.~Yamamoto,
  %``Photonic chiral vortical effect,''
  Phys.\ Rev.\ D {\bf 96}, no. 5, 051902 (2017)
  doi:10.1103/PhysRevD.96.051902
  [arXiv:1702.08886 [hep-th]].
  %%CITATION = doi:10.1103/PhysRevD.96.051902;%%
  %3 citations counted in INSPIRE as of 11 Dec 2017

\bibitem{Charbonneau:2009ax} 
  J.~Charbonneau and A.~Zhitnitsky,
  %``Topological Currents in Neutron Stars: Kicks, Precession, Toroidal Fields, and Magnetic Helicity,''
  JCAP {\bf 1008}, 010 (2010).


%\cite{Dvornikov:2016gdo}
\bibitem{Dvornikov:2016gdo} 
  M.~Dvornikov and V.~B.~Semikoz,
  %``Non-conservation of the neutrino current in a hot plasma of the early universe,''
  arXiv:1603.07946 [astro-ph.CO].
  %%CITATION = ARXIV:1603.07946;%%

\bibitem{Kaplan:2016drz} 
  D.~B.~Kaplan, S.~Reddy and S.~Sen,
  %``Energy Conservation and the Chiral Magnetic Effect,''
  arXiv:1612.00032 [hep-ph].


  
  %\cite{Mueller:2017lzw}
\bibitem{Mueller:2017lzw} 
  N.~Mueller and R.~Venugopalan,
  %``The chiral anomaly, Berry's phase and chiral kinetic theory, from world-lines in quantum field theory,''
  arXiv:1701.03331 [hep-ph].
  %%CITATION = ARXIV:1701.03331;%%
  %7 citations counted in INSPIRE as of 05 Dec 2017
  
%\cite{long-paper}
\bibitem{Mueller:2017arw}
N.~Mueller and R.~Venugopalan,
  %``Worldline construction of a covariant chiral kinetic theory,''
  Phys.\ Rev.\ D {\bf 96}, no. 1, 016023 (2017). 
  
  
%\cite{Polyakov:1987ez}
\bibitem{Polyakov:1987ez} 
  A.~M.~Polyakov,
  %``Gauge Fields and Strings,''
  Contemp.\ Concepts Phys.\  {\bf 3}, 1 (1987).
  %%CITATION = 00380,3,1;%%
  %110 citations counted in INSPIRE as of 05 Dec 2017
  
  
 
  
%\cite{Strassler:1992zr}
\bibitem{Strassler:1992zr} 
  M.~J.~Strassler,
  %``Field theory without Feynman diagrams: One loop effective actions,''
  Nucl.\ Phys.\ B {\bf 385}, 145 (1992)
  %doi:10.1016/0550-3213(92)90098-V
  [hep-ph/9205205].
  %%CITATION = %doi:10.1016/0550-3213(92)90098-V;%%
  %336 citations counted in INSPIRE as of 06 Dec 2016

\bibitem{Mondragon:1995va} 
  M.~Mondragon, L.~Nellen, M.~G.~Schmidt and C.~Schubert,
  %``Yukawa couplings for the spinning particle and the worldline formalism,''
  Phys.\ Lett.\ B {\bf 351}, 200 (1995)
  %doi:10.1016/0370-2693(95)00337-K
  [hep-th/9502125].

%\cite{Mondragon:1995ab}
\bibitem{Mondragon:1995ab} 
  M.~Mondragon, L.~Nellen, M.~G.~Schmidt and C.~Schubert,
  %``Axial couplings on the worldline,''
  Phys.\ Lett.\ B {\bf 366}, 212 (1996)
  %doi:10.1016/0370-2693(95)01392-X
  [hep-th/9510036].
  %%CITATION = %doi:10.1016/0370-2693(95)01392-X;%%
  %35 citations counted in INSPIRE as of 19 Jan 2017
  
  
%\cite{JalilianMarian:1999xt}
\bibitem{JalilianMarian:1999xt} 
  J.~Jalilian-Marian, S.~Jeon, R.~Venugopalan and J.~Wirstam,
  %``Minding one's P's and Q's: From the one loop effective action in quantum field theory to classical transport theory,''
  Phys.\ Rev.\ D {\bf 62}, 045020 (2000)
  %doi:10.1103/PhysRevD.62.045020
  [hep-ph/9910299].
  %%CITATION = %doi:10.1103/PhysRevD.62.045020;%%
  %16 citations counted in INSPIRE as of 06 Dec 2016
  
  
%\cite{Schubert:2001he}
\bibitem{Schubert:2001he} 
  C.~Schubert,
  %``Perturbative quantum field theory in the string inspired formalism,''
  Phys.\ Rept.\  {\bf 355}, 73 (2001)
  %doi:10.1016/S0370-1573(01)00013-8
  [hep-th/0101036].
  %%CITATION = %doi:10.1016/S0370-1573(01)00013-8;%%
  %222 citations counted in INSPIRE as of 06 Dec 2016
  
  
  
 %\cite{Bastianelli:2006rx}
\bibitem{Bastianelli:2006rx} 
  F.~Bastianelli and P.~van Nieuwenhuizen, {\it Path integrals and anomalies in curved space}, Cambridge University Press (2006).
  %%CITATION = INSPIRE-728108;%%
  %12 citations counted in INSPIRE as of 06 Dec 2016 

%\cite{Hernandez:2008db}
\bibitem{Hernandez:2008db} 
  A.~Hernandez, T.~Konstandin and M.~G.~Schmidt,
  %``Sizable CP Violation in the Bosonized Standard Model,''
  Nucl.\ Phys.\ B {\bf 812}, 290 (2009)
  %doi:10.1016/j.nuclphysb.2008.12.021
  [arXiv:0810.4092 [hep-ph]].
  %%CITATION = %doi:10.1016/j.nuclphysb.2008.12.021;%%
  %26 citations counted in INSPIRE as of 19 Jan 2017
  
  
%\cite{Corradini:2015tik}
\bibitem{Corradini:2015tik} 
  O.~Corradini and C.~Schubert,
  %``Spinning Particles in Quantum Mechanics and Quantum Field Theory,''
  arXiv:1512.08694 [hep-th].
  %%CITATION = ARXIV:1512.08694;%%
  
  
  
%\cite{Berezin:1976eg}
\bibitem{Berezin:1976eg} 
  F.~A.~Berezin and M.~S.~Marinov,
  %``Particle Spin Dynamics as the Grassmann Variant of Classical Mechanics,''
  Annals Phys.\  {\bf 104}, 336 (1977).
  %doi:10.1016/0003-4916(77)90335-9
  %%CITATION = %doi:10.1016/0003-4916(77)90335-9;%%
  %583 citations counted in INSPIRE as of 06 Dec 2016 
  
%\cite{Ohnuki:1978jv,Balachandran:1976ya,Balachandran:1977ub,Barducci:1976xq, Barducci:1982yw,Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Ohnuki:1978jv} 
  Y.~Ohnuki and T.~Kashiwa,
  %``Coherent States of Fermi Operators and the Path Integral,''
  Prog.\ Theor.\ Phys.\  {\bf 60}, 548 (1978).
  %doi:10.1143/PTP.60.548
  %%CITATION = %doi:10.1143/PTP.60.548;%%
  %97 citations counted in INSPIRE as of 06 Dec 2016


%\cite{Balachandran:1976ya,Balachandran:1977ub,Barducci:1976xq, Barducci:1982yw,Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Balachandran:1976ya} 
  A.~P.~Balachandran, P.~Salomonson, B.~S.~Skagerstam and J.~O.~Winnberg,
  %``Classical Description of Particle Interacting with Nonabelian Gauge Field,''
  Phys.\ Rev.\ D {\bf 15}, 2308 (1977).
  %doi:10.1103/PhysRevD.15.2308
  %%CITATION = %doi:10.1103/PhysRevD.15.2308;%%
  %147 citations counted in INSPIRE as of 06 Dec 2016  
  
%\cite{Balachandran:1977ub,Barducci:1976xq, Barducci:1982yw,Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Balachandran:1977ub} 
  A.~P.~Balachandran, S.~Borchardt and A.~Stern,
  %``Lagrangian and Hamiltonian Descriptions of Yang-Mills Particles,''
  Phys.\ Rev.\ D {\bf 17}, 3247 (1978).
  %doi:10.1103/PhysRevD.17.3247
  %%CITATION = %doi:10.1103/PhysRevD.17.3247;%%
  %84 citations counted in INSPIRE as of 06 Dec 2016

\bibitem{Barducci:1976xq} 
  A.~Barducci, R.~Casalbuoni and L.~Lusanna,
  %``Classical Scalar and Spinning Particles Interacting with External Yang-Mills Fields,''
  Nucl.\ Phys.\ B {\bf 124}, 93 (1977).
  
  
 %\cite{Galvao:1980cu}
\bibitem{Galvao:1980cu} 
  C.~A.~P.~Galvao and C.~Teitelboim,
  %``Classical Supersymmetric Particles,''
  J.\ Math.\ Phys.\  {\bf 21}, 1863 (1980).
  doi:10.1063/1.524603
  %%CITATION = doi:10.1063/1.524603;%%
  %107 citations counted in INSPIRE as of 11 Dec 2017

  
%\cite{Barducci:1982yw,Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Barducci:1982yw} 
  A.~Barducci,
  %``Pseudoclassical Description of Relativistic Spinning Particles With Anomalous Magnetic Moment,''
  Phys.\ Lett.\  {\bf 118B}, 112 (1982).
  %doi:10.1016/0370-2693(82)90611-6
  %%CITATION = %doi:10.1016/0370-2693(82)90611-6;%%
  %12 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Brink:1976uf} 
  L.~Brink, P.~Di Vecchia and P.~S.~Howe,
  %``A Lagrangian Formulation of the Classical and Quantum Dynamics of Spinning Particles,''
  Nucl.\ Phys.\ B {\bf 118}, 76 (1977).
  %doi:10.1016/0550-3213(77)90364-9
  %%CITATION = %doi:10.1016/0550-3213(77)90364-9;%%
  %417 citations counted in INSPIRE as of 06 Dec 2016
  
  %\cite{Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn}
\bibitem{Gitman:1996wk} 
  D.~M.~Gitman,
  %``Path integrals and pseudoclassical description for spinning particles in arbitrary dimensions,''
  Nucl.\ Phys.\ B {\bf 488}, 490 (1997)
  %doi:10.1016/S0550-3213(96)00691-8
  [hep-th/9608180].
  %%CITATION = %doi:10.1016/S0550-3213(96)00691-8;%%
  %23 citations counted in INSPIRE as of 20 Apr 2017
  
  %\cite{Gavrilov:2001hk}
\bibitem{Gavrilov:2001hk} 
  S.~P.~Gavrilov and D.~M.~Gitman,
  %``Quantization of spinning particle in arbitrary background,''
  Class.\ Quant.\ Grav.\  {\bf 18}, 2989 (2001)
  %doi:10.1088/0264-9381/18/15/312
  [hep-th/0104121].
  %%CITATION = %doi:10.1088/0264-9381/18/15/312;%%
  %4 citations counted in INSPIRE as of 20 Apr 2017
  
  %\cite{Naka:1979pn}
\bibitem{Naka:1979pn} 
  S.~Naka,
  %``Canonical Approach To Spinning Particles,''
  Prog.\ Theor.\ Phys.\  {\bf 62}, 278 (1979).
  %doi:10.1143/PTP.62.278
  %%CITATION = %doi:10.1143/PTP.62.278;%%
  %1 citations counted in INSPIRE as of 20 Apr 2017

  
  
%\cite{AlvarezGaume:1983ig}
\bibitem{AlvarezGaume:1983ig} 
  L.~Alvarez-Gaume and E.~Witten,
  %``Gravitational Anomalies,''
  Nucl.\ Phys.\ B {\bf 234}, 269 (1984).
  %doi:10.1016/0550-3213(84)90066-X
  %%CITATION = %doi:10.1016/0550-3213(84)90066-X;%%
  %1389 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{AlvarezGaume:1983at}
\bibitem{AlvarezGaume:1983at} 
  L.~Alvarez-Gaume,
  %``Supersymmetry and the Atiyah-Singer Index Theorem,''
  Commun.\ Math.\ Phys.\  {\bf 90}, 161 (1983).
  %doi:10.1007/BF01205500
  %%CITATION = %doi:10.1007/BF01205500;%%
  %299 citations counted in INSPIRE as of 06 Dec 2016


%\cite{D'Hoker:1995ax,D'Hoker:1995bj}
\bibitem{D'Hoker:1995ax} 
  E.~D'Hoker and D.~G.~Gagne,
  %``Worldline path integrals for fermions with scalar, pseudoscalar and vector couplings,''
  Nucl.\ Phys.\ B {\bf 467}, 272 (1996)
  %doi:10.1016/0550-3213(96)00125-3
  [hep-th/9508131].
  %%CITATION = %doi:10.1016/0550-3213(96)00125-3;%%
  %63 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{D'Hoker:1995bj}
\bibitem{D'Hoker:1995bj} 
  E.~D'Hoker and D.~G.~Gagne,
  %``Worldline path integrals for fermions with general couplings,''
  Nucl.\ Phys.\ B {\bf 467}, 297 (1996)
  %doi:10.1016/0550-3213(96)00126-5
  [hep-th/9512080].
  %%CITATION = %doi:10.1016/0550-3213(96)00126-5;%%
  %51 citations counted in INSPIRE as of 06 Dec 2016
  


%\cite{Deguchi:2005pc,Fujikawa:2005tv,Fujikawa:2005cn}
\bibitem{Deguchi:2005pc} 
  S.~Deguchi and K.~Fujikawa,
  %``Second quantized formulation of geometric phases,''
  Phys.\ Rev.\ A {\bf 72}, 012111 (2005)
  %doi:10.1103/PhysRevA.72.012111
  [hep-th/0501166].
  %%CITATION = %doi:10.1103/PhysRevA.72.012111;%%
  %11 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Fujikawa:2005tv,Fujikawa:2005cn}
\bibitem{Fujikawa:2005tv} 
  K.~Fujikawa,
  %``Geometric phases and hidden local gauge symmetry,''
  Phys.\ Rev.\ D {\bf 72}, 025009 (2005)
  %doi:10.1103/PhysRevD.72.025009
  [hep-th/0505087].
  %%CITATION = %doi:10.1103/PhysRevD.72.025009;%%
  %12 citations counted in INSPIRE as of 06 Dec 2016
  
%\cite{Fujikawa:2005cn}
\bibitem{Fujikawa:2005cn} 
  K.~Fujikawa,
  %``Quantum anomaly and geometric phase: Their basic differences,''
  Phys.\ Rev.\ D {\bf 73}, 025017 (2006)
  %doi:10.1103/PhysRevD.73.025017
  [hep-th/0511142].
  %%CITATION = %doi:10.1103/PhysRevD.73.025017;%%
  %5 citations counted in INSPIRE as of 06 Dec 2016
  
 
%\cite{Fujikawa:2017ych}
\bibitem{Fujikawa:2017ych} 
  K.~Fujikawa,
  %``Characteristics of Chiral Anomaly in View of Various Applications,''
  arXiv:1709.08181 [hep-th].
  %%CITATION = ARXIV:1709.08181;%%
  
%\cite{Bargmann:1959gz,Wong:1970fu} 
\bibitem{Bargmann:1959gz} 
  V.~Bargmann, L.~Michel and V.~L.~Telegdi,
  %``Precession of the polarization of particles moving in a homogeneous electromagnetic field,''
  Phys.\ Rev.\ Lett.\  {\bf 2}, 435 (1959).
  %doi:10.1103/PhysRevLett.2.435
  %%CITATION = %doi:10.1103/PhysRevLett.2.435;%%
  %582 citations counted in INSPIRE as of 22 Nov 2016
  
%\cite{Wong:1970fu}
\bibitem{Wong:1970fu} 
  S.~K.~Wong,
  %``Field and particle equations for the classical Yang-Mills field and particles with isotopic spin,''
  Nuovo Cim.\ A {\bf 65}, 689 (1970).
  %doi:10.1007/BF02892134
  %%CITATION = %doi:10.1007/BF02892134;%%
  %377 citations counted in INSPIRE as of 27 Nov 2016



\bibitem{BBGKY}
N. N. Bogoliubov, JETP 16 (8) 1946, 691-702;
J. Yvon, Actual. Sci. \& Indust. 203 (1937);
J.G. Kirkwood, Journal of Chemical Physics 14 (3), 180

\bibitem{deGroot}
S.~de~Groot, W.~.A.~Leeuwen, C.~G.~van~Weert,
 ``Relativistic Kinetic Theory'', North-Holland Pub. Co. (1980)

\bibitem{Landau}
	V.~B.~Berestetskii, E.~M.~Lifshitz and L.~P.~Pitaevskii,
  ``Quantum Electrodynamics'', Butterworth-Heinemann  (1982)
  %%CITATION = INSPIRE-185062;%%
  %15 citations counted in INSPIRE as of 11 Dec 2017
 
  
\end{thebibliography}

\documentclass{PoS}
\usepackage{graphicx,subfigure}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{bm}
\usepackage{color}
\usepackage{comment}
\usepackage{slashed}

\def \be {\begin{equation} }
\def \ee {\end{equation}}

\def \bem {\begin{multline}}
\def \eem {\end{multline}}

\def \bes {\begin{subequations} }
\def \ees {\end{subequations}}

\newcommand{\Eq}[1]{Eq.~(\ref{#1})}

\title{World-line formulation of chiral kinetic theory in topological background gauge fields}

\ShortTitle{World-line approach to chiral kinetic theory}

\author{Niklas Mueller\\
        Physics Department, Brookhaven National Laboratory, Bldg. 510A, Upton, NY 11973, USA\\
        E-mail: \email{nmueller@bnl.gov}}

\author{\speaker{Raju Venugopalan}\\
        Physics Department, Brookhaven National Laboratory, Bldg. 510A, Upton, NY 11973, USA\\
        E-mail: \email{raju@bnl.gov}}

\abstract{In heavy-ion collisions, an interesting question of phenomenological relevance is how the chiral imbalance generated at early times persists through a fluctuating background of sphalerons in addition to other "non-anomalous" interactions with the QGP. To address this question, we construct a relativistic chiral kinetic theory using the world-line formulation of quantum field theory. We outline how Berry's phase arises in this framework, and how its effects can be clearly distinguished from those arising from the chiral anomaly. We further outline how this framework can be matched to classical statistical simulations at early times and to anomalous chiral hydrodynamics at late times.}

\FullConference{Critical Point and Onset of Deconfinement - CPOD2017\\
		7-11 August, 2017\\
		The Wang Center, Stony Brook University, Stony Brook, NY}

\begin{document}

\section{Introduction}
Experimental searches for messengers of CP- and P-odd phenomena in ultra-relativistic heavy ion collisions have aroused much interest and are a prime motivation for significant theoretical effort in this area in recent years. Topological sphaleron transitions~\cite{Klinkhamer:1984di,Dashen:1974ck,Soni:1980ps,Boguta:1983xs,Forgacs:1983yu} were first conjectured to exist in the context of electroweak baryogenesis where they were inferred to be responsible for the matter-antimatter asymmetry in the universe~\cite{Sakharov:1967dj,Riotto:1999yt,Cohen:1993nk,Rubakov:1996vz}. In QCD, these topological 
transitions generate a Chiral Magnetic Effect~\cite{Kharzeev:2007jp,Fukushima:2008xe,Kharzeev:2015znc} in heavy-ion collisions, whereby a combination of CP- and P-odd effects and
strong external (Abelian) magnetic fields can lead to  correlations between charged particles~\cite{Kharzeev:2015znc,Skokov:2016yrj}. However the short lifetime of the external magnetic fields \cite{Skokov:2009qp,Deng:2012pc} suggests that  the CME is largest during the early stages of a collision where the system is far off-equilibrium. The matter at these early times is a strongly correlated "Glasma"-state~\cite{Lappi:2006fp,Gelis:2006dv,Blaizot:2016qgz} formed from the decay of highy occupied Color Glass Condensates~\cite{McLerran:1993ni,McLerran:1993ka,Gelis:2010nm} and its dynamics is captured by ab initio classical-statistical real time lattice simulations ~\cite{Berges:2013eia,Berges:2013fga,Gelis:2016upa}.  These methods can be extended to compute off-equilibrium sphaleron rates in the Glasma~\cite{Mace:2016svc}, which are seen to be significantly larger than the rate of sphaleron transitions in the thermalized Quark-Gluon Plasma \cite{Moore:2010jd}. 

First simulations of dynamical fermions in the gauge fields generating these sphaleron transitions show unambiguously that a CME develops in the presence of the external magnetic field~\cite{Mueller:2016ven,Mace:2016shq,Mace:2017wcl}.
The classical-statistical framework breaks down when occupancies decrease below unity due to the expansion of the overoccupied Glasma.  In this
dilute regime, the results of real-time lattice simulations must be understood as providing the initial conditions for kinetic descriptions
for a weakly coupled dynamical fluid of quasi-particles. Following the space-time evolution of the CME therefore requires that one develops a Boltzmann transport theory for
relativistic chiral fermions which also accounts for its interactions with the topological transitions in the evolving fluid. Such an {\it ab initio} framework is of the utmost importance for systematic analyses of the CME phenomenology, since it can be matched at later times to (anomalous) hydrodynamics ~\cite{Son:2009tf,Nair:2011mk,Gursoy:2014aka,Hongo:2013cqa,Karabali:2014vla,Hirono:2014oda,
Yin:2015fca}.  Several suggestions for a chiral kinetic theory have been made, utilizing models of point particles with
a Berry monopole~\cite{Berry:1984jv}; the latter has been shown to arise in an adiabatic limit for either massive nonrelativistic particles or for massless relativistic particles with a large chemical potential~\cite{Son:2012wh,Stephanov:2012ki,Son:2012zy,Chen:2013iga,Chen:2014cla,Stone:2013sga,Dwivedi:2013dea,Manuel:2014dza,Stone:2014fja,Manuel:2015zpa,Chen:2015gta,Sun:2016nig,Hidaka:2016yjf}.
This framework can be extended to construct an "anomalous B\"odeker theory" \cite{Bodeker:1999ey,Litim:1999id}, which can then be matched to classical-statistical simulations at early times in heavy-ion collisions and to anomalous hydrodynamics at late times. 
It is not clear that early time dynamics in heavy-ion collisions respects these asymptotics and that the inclusion of a Berry term in kinetic descriptions is justified in this context. On the other hand, the physics of the chiral anomaly is ubiquitous and it is essential that chiral kinetic theory frameworks incorporate this physics. An important development is due to \cite{Akamatsu:2013pjd,Akamatsu:2014yza,Yamamoto:2015gzz,Akamatsu:2015kau,Yamamoto:2017uul}, highlighting the possible applicability of such frameworks to astrophysical situations~\cite{Charbonneau:2009ax,Dvornikov:2016gdo,Kaplan:2016drz}. 

In this talk, we will  give an overview of our recent work \cite{Mueller:2017lzw,Mueller:2017arw} in developing a consistent Lorentz covariant transport theory within the world-line 
approach to quantum field theory~\cite{Polyakov:1987ez,Strassler:1992zr,Mondragon:1995va,Mondragon:1995ab,JalilianMarian:1999xt,Schubert:2001he,Bastianelli:2006rx,Hernandez:2008db,Corradini:2015tik}. 
A central ingredient, going back to seminal work by Berezin and Marinov \cite{Berezin:1976eg}
and related contemporaneous work~\cite{Ohnuki:1978jv,Balachandran:1976ya,Balachandran:1977ub,Barducci:1976xq, Galvao:1980cu,Barducci:1982yw,Brink:1976uf, Gitman:1996wk,Gavrilov:2001hk,Naka:1979pn} is the description of spin, color and other internal symmetries via Grassmann variables. The outline of the manuscript is as follow.  In section~\ref{sec:introWL}, we will give a short overview over the world-line representation of the one-loop effective action for fermions. Subsequently,  we
connect the emergence of the chiral anomaly to the existence of fermionic zero modes in the world-line path integral. In section \ref{sec:adiabatic}, we show systematically how Berry's phase arises in our framework. Our explicit computation clearly demonstrates that the topology of the chiral anomaly
and that of Berry's phase  are distinct and should not be confused with each other. For instance, the latter  vanishes when we relax the adiabaticity assumption while the former is robust. In section \ref{sec:CKT}, we outline the elements of a many-body phase space description for relativistic fermions, wherein spin is represented by anti-commuting Grassman coordinates. 

\section{World-line formalism}\label{sec:introWL}
We will briefly here review the world-line representation of quantum field theory
and refer the reader to \cite{Mueller:2017lzw,Mueller:2017arw} for more details.  
We shall restrict ourselves to QED; the extension to QCD
is straightforward. The Euclidean action for massless Dirac fermions in a background vector field $A$ and an auxilliary axial-vector field $B$ is 
\begin{align}
S[A,B]=\int d^4x\;\bar{\psi}\left(i\slashed{\partial}+\slashed{A}+\gamma_5\slashed{B} \right)\psi\,,\label{eq:classicalaction}
\end{align}
The fermion one-loop effective action is given by the fermion determinant $-W[A,B]=\log\det(\theta)$, which can be split 
into a real and and an imaginary part, the latter corresponding to the phase of the determinant. This phase is well known to be responsible for the chiral anomaly \cite{AlvarezGaume:1983ig,AlvarezGaume:1983at}. Remarkably, one can obtain heat-kernel expressions for {\it both} the real and the imaginary part
of the effective action in the world-line formalism~ \cite{D'Hoker:1995ax,D'Hoker:1995bj}, which are expressed through quantum mechanical path integrals of a set of bosonic ($x^\mu$) and anticommuting Grassmann variables $(\psi^\mu,\psi^5,\psi^6)$ quantized on a closed loop.  The real part of the effective action in this formalism can be expressed as 
\begin{align}
W_\mathbb{R}=\frac{1}{8}\int\limits_0^\infty\frac{dT}{T}\mathcal{N}\int\limits_{P}
\mathcal{D}x\int\limits_{AP}\mathcal{D}\psi\;\text{tr}\exp{\Big\{-\int\limits_0^Td\tau\;\mathcal{L}(\tau)\Big\}}\label{eq:realpart}.
\end{align}
where the point particle Lagrangian, $\mathcal{L}=\text{diag}(\mathcal{L}_{L},\mathcal{L}_{R})$,  is given by 
\begin{align}\label{eq:lagrangianexplicit}
&\mathcal{L}_{L/R}=\frac{\dot{x}^2}{2\epsilon}+\frac{1}{2}\psi_a\dot{\psi}_a -i\dot{x}_\mu (A\pm B)_\mu-\frac{i\mathcal{E}}{2}\psi_\mu\psi_\nu F_{\mu\nu}[A\pm B]\,,
\end{align}
with $a=1,\cdots,6$. Here $\epsilon$ is an einbein parameter related to the reparametrization invariance of the world-line. The imaginary part of the effective action can likewise be expressed as a path integral with the same point particle Lagrangian, 
but with a few key differences. Firstly, the imaginary part of the effective action contains a trace insertion
which turns the anti-periodic boundary conditions for the Grassmann variables $\psi$ into periodic ones, thus
being responsible for the existence of fermionic zero modes. Further, this part of the effective action contains an integral over a variable $\alpha$ which has the limits $\pm 1$.  The point particle Lagrangian is identical to that in \Eq{eq:lagrangianexplicit} except that $B\rightarrow \alpha B$. For $\alpha\neq \pm 1$, the imaginary part of the effective action violates chiral symmetry explicitly. The anomaly relation
can be computed directly by varying the imaginary part of the effective action with respect to $B$ and subsequently setting $B=0$: 
\begin{align}\label{eq:finalanomaly}
\partial_\mu \langle j^5_\mu(y)\rangle\equiv
\partial_\mu\frac{i\delta W_\mathbb{I}}{\delta B_\mu(y)}\Big|_{B=0}=-\frac{1}{16\pi^2}\epsilon^{\mu\nu\rho\sigma}F_{\mu\nu}(y)F_{\rho\sigma}(y)\, .
\end{align}
Details of these derivations are worked out in \cite{Mueller:2017lzw,Mueller:2017arw}. Our result clearly shows that the chiral anomaly arises from fermionic zero modes, that uniquely contribute to the imaginary part of the effective action and are absent in the real part of the effective action.

\section{Berry's phase from the real part of the fermion effective action}\label{sec:adiabatic}
In this section, we will investigate the consequences of taking the limit of large masses or large chemical potential in the real part of the effective action and show how Berry's phase further arises in an adiabatic approximation. This cleanly illustrates that the topology of Berry's phase is unrelated to that of the anomaly, as argued previously by Fujikawa and collaborators in more specialized contexts~\cite{Deguchi:2005pc,Fujikawa:2005tv,Fujikawa:2005cn}.

We will first continue the world-line Lagrangian in \Eq{eq:lagrangianexplicit} to Minkowskian metric $g=\text{diag}(-,+,+,+)$. 
To ensure that excitations follow the correct spectrum dictated by the Dirac equation, we introduce the so-called helicity
constraint $\frac{\dot{x}_\mu\psi^\mu}{2\mathcal{E}}+m\psi_5=0$ and a corresponding anti-commuting Lagrange multiplier. Generalizing our discussion to the case of
a massive fermion, and assuming proper time gauge, $\tau=ct\sqrt{1-(\mathbf{v}/c)^2}$, $\sqrt{-\dot{z}^2}=1$, where $\dot{z}^2=(dx^\mu/d\tau)^2$, we find~\cite{Mueller:2017arw} 
\begin{align}
\label{eq:NRLagrangean1}
&\mathcal{L}= -\frac{m_R \,c \,z}{2}\left( 1+\frac{m^2}{m_R^2}\right)+\frac{i}{2}\left(\boldsymbol{\psi} \dot{\boldsymbol{\psi}}- \psi_0\dot{\psi}_0 \right) 
+\frac{\dot{x}_\mu A^\mu(x)}{c} - \frac{iz}{m_R\,c}\,\psi^0 F_{0i}\psi^i- \frac{iz}{2m_R\,c}\,\psi^i F_{ij}\psi^j \, ,
\end{align}
where $m_R^2=m^2+i\psi\cdot F\cdot \psi$. To take the nonrelativistic limit, we expand the action $S=\int d\tau \mathcal{L}=\int dt\,c\sqrt{1-(\mathbf{v}/c)^2} \mathcal{L}$ in powers of
$(v/c)^2$, to obtain 
\begin{align}
\label{eq:NRLagrangean2}
\mathcal{L}_{NR}&=-mc^2+\frac{1}{2}m \boldsymbol{v}^2+\frac{i}{2}\left(\boldsymbol{\psi} \dot{\boldsymbol{\psi}}- \psi_0\dot{\psi}_0 \right) - A^0 +\frac{\boldsymbol{v}}{c}\cdot\boldsymbol{A}
+\frac{ \boldsymbol{S}\cdot(\left[ {\boldsymbol{v}/c}-{\boldsymbol{A}/(mc^2)}\right]\times\boldsymbol{E})}{mc}+\frac{\boldsymbol{S}\cdot\boldsymbol{B}}{m} + O\Bigg(\frac{v^3}{c^3}\Bigg)\,.
\end{align}
The corresponding Hamiltonian is familiar to us from atomic physics: 
\begin{align}\label{eq:NRhamilt}
H\equiv mc^2&+\frac{\left(\boldsymbol{p}-\frac{\boldsymbol{A}}{c}\right)^2}{2m}+A^0(x)
-\frac{ \boldsymbol{S}\cdot(\left[ {\boldsymbol{v}/c}-{\boldsymbol{A}/(mc^2)}\right]\times\boldsymbol{E})}{2mc}-\frac{\boldsymbol{B}\cdot\boldsymbol{S}}{m}\,.
\end{align}
A similar expression is found for the case of a large chemical potential $\mu \gg m$, where the chemical potential 
takes over the role of the mass in \Eq{eq:NRhamilt}. The world-line path integral for the \textit{real} part of the fermion
effective action in the nonrelativistic adiabatic limit can be simplified to read as
\begin{align}\label{eq:Berrypath}
W_\mathbb{R}=\int \mathcal{D}x\mathcal{D}p\;\exp{\Big(i \int dt\;\Big[\dot{\mathbf{x}}\cdot\mathbf{p}-\tilde {H}\Big] \Big)} \,,
\end{align}
with ${\tilde H} = mc^2+\frac{(\boldsymbol{p}-\boldsymbol{A}/c)^2}{2m}+A^0(x)-\dot{\mathbf{p}}\cdot\boldsymbol{\mathcal{A}}(\boldsymbol{p})$. The last term includes the well known Berry phase $\boldsymbol{\mathcal{A}}(\boldsymbol{p})$ $\equiv -i \langle \psi^+(\boldsymbol{p})|\boldsymbol{\nabla}_p| \psi^+(\boldsymbol{p})\rangle$. In deriving this expression, the adiabatic assumption consisted of assuming that the particle's spin was slaved to the direction of the magnetic field and that spin flips are suppressed. Comparing \Eq{eq:Berrypath} to the results of \cite{Son:2012wh,Stephanov:2012ki,Son:2012zy,Chen:2013iga},
we should note a few points: While the emergence of the anomaly was tied to the existence of a Berry phase in \cite{Son:2012wh,Stephanov:2012ki,Son:2012zy,Chen:2013iga}, we see here that we can recover the very same Berry monopole from the
real part of the effective action while the anomaly is tied to the imaginary part.  As stated earlier, our results agree with observations
made by Fujikawa and collaborators in more specific contexts \cite{Deguchi:2005pc,Fujikawa:2005tv,Fujikawa:2005cn,Fujikawa:2017ych}. 
\section{Towards Chiral Kinetic Theory}\label{sec:CKT}
The world-line approach has been employed previously~\cite{JalilianMarian:1999xt} for spinless colored particles to derive the B\"odeker effective kinetic theory for finite temperature QCD~\cite{Bodeker:1999ey,Litim:1999id}. Starting from the Schwinger-Keldysh formulation of the world-line approach,  a quasi-classical many-body description of point particles was obtained by taking the saddle point of the world-line action. The initial density
matrix in the Schwinger-Keldysh (SK) formulation thus represents a stochastic ensemble of classical phase space configurations. For the case of spin, the first step is to identify the 
extended semi-classical phase space to include spinning degrees of freedom; this is analogous to the case of colored particles where the phase space was extended to include their color charge. The natural variables in the spinning case are the Grassmannian variables $\psi$. This gives us  $(x^\mu,P^\mu)\rightarrow(x^\mu,P^\mu,\psi^\mu,\psi_5)$; as discussed previously in \cite{Mueller:2017arw}, $\psi_6$ is not dynamical and is therefore dropped.

The dynamics underlying the microscopic phase space distribution for $N$ particles $f(x,P,\psi)$ 
is governed by the many-body Hamiltonian,
\begin{align}\label{eq:MassiveHamiltonian}
H=\frac{\epsilon}{2}\left(P^2 + m^2 + i \psi^\mu F_{\mu\nu}\psi^\nu \right) + \frac{i}{2}\left(P_\mu \psi^\mu + m\psi_5 \right)\chi
\end{align}
where $P^\mu\equiv p^\mu-A^\mu$ and summation over all particles is implied. In this expression, $\epsilon$ is a commuting Lagrange
multiplier ("einbein") enforcing the mass shell condition while $\chi$ is an anti-commuting Lagrange
multiplier related to the Dirac equation. Both represent gauge parameters with respect to first class constraints and must be fixed before extracting 
physical results. The dynamics of the system can be understood from
the Liouville equation~\cite{Mueller:2017lzw}
\begin{align}\label{eq:Liouville}
0=\{f,H \}=f\Big( \frac{\overleftarrow{\partial}}{\partial x^\mu}\dot{x}^\mu +\frac{\overleftarrow{\partial}}{\partial P^\mu}\dot{P}^\mu +\frac{\overleftarrow{\partial}}{\partial \psi^\mu}\dot{\psi}^\mu +\frac{\overleftarrow{\partial}}{\partial \psi_5}\dot{\psi}_5\Big)\,,
\end{align}
where
\begin{align}
\dot{P}^\mu & =\epsilon P_\alpha F^{\mu\alpha}- \frac{i\epsilon}{2}\psi^\alpha \partial^\mu F_{\alpha\beta}\psi^\beta + \frac{i}{2} F^{\mu\alpha} \psi_\alpha\,\chi\,,\\
\dot{x}^\mu &= \epsilon P^\mu + \frac{i}{2}\psi^\mu \,\chi\,,\\
\dot{\psi}^\mu &= \epsilon F^{\mu\alpha}\psi_\alpha + \frac{P^\mu}{2}\chi\,,\\
\dot{\psi}_5 & = \frac{m}{2}\chi\,.
\end{align}
These equations of motion are generalizations~\cite{Barducci:1976xq}  of the Bargmann-Michel-Telegdi (BMT) equations \cite{Bargmann:1959gz} for spinning particles. For QCD, their equivalents are the Wong equations~\cite{Wong:1970fu}.

Important conceptual issues must be addressed in order to turn this Liouville equation into a transport theory
for fermions whereby both effects of the chiral anomaly and of particle scattering are accounted for. Firstly, since it is impractical to follow every single phase space trajectory via \Eq{eq:Liouville}, a stochastic approach must be
developed that converts the microscopic phase space distribution $f$ into a macroscopic one-particle {\it probability distribution} $\bar{f}$.
Such an analysis should yield the scattering terms in the generalized Boltzmann equation for $\bar{f}$.
The Wigner transform of the initial density matrix in the SK path integral represents a Gibbs ensemble of phase space distributions
in the saddle point limit. This motivates the definition of the distribution functions as
\begin{align}
f\equiv \langle f \rangle + \delta f= \bar{f}+ \delta f
\end{align} 
and similarly for the gauge fields within the SK path integral:
\begin{align}
A^\mu\equiv \langle A^\mu \rangle + \delta A^\mu= \bar{A}^\mu+ \delta A^\mu\,.
\end{align} 
For QED, this also implies $F^{\mu\nu}=\bar{F}^{\mu\nu}+\delta F^{\mu\nu}$. Taking into consideration the fact that the Liouville equation is 
 embedded in the SK path integral, it can be split into separate contributions for the background and fluctuations. The former is given by
\begin{align}
\bar{f}\Big(\frac{\overleftarrow{\partial}}{\partial x^\mu}&\Big[ 
\epsilon P^\mu + \frac{i}{2}\psi^\mu \chi\Big] + \frac{\overleftarrow{\partial}}{\partial P^\mu}\Big[  \epsilon \bar{F}^{\mu\alpha} P_\alpha - \frac{i \epsilon}{2}\psi^\alpha \partial^\mu \bar{F}_{\alpha\beta} \psi^\beta \nonumber\\&+ \frac{i}{2}\bar{F}^{\mu\alpha}\psi_\alpha \chi\Big]
+\frac{\overleftarrow{\partial}}{\partial \psi^\mu}\Big[ 
\epsilon \bar{F}^{\mu\alpha}\psi_\alpha + \frac{P^\mu}{2}\chi
\Big]+\frac{\overleftarrow{\partial}}{\partial \psi_5}\Big[ \frac{m}{2}\chi\Big]\Big)=C[\delta f, \delta F]\,,
\label{Eq:lioudetails2}
\end{align}
where
\begin{align}
C[\delta f, \delta F]\equiv -\epsilon \langle \delta f\frac{\overleftarrow{\partial}}{\partial P^\mu}\,\delta F^{\mu\alpha} \rangle P_\alpha -
\epsilon \langle \delta f\frac{\overleftarrow{\partial}}{\partial \psi^\mu}\,\delta F^{\mu\alpha} \rangle & \psi_\alpha+\frac{i \epsilon}{2} \langle \delta f\frac{\overleftarrow{\partial}}{\partial P^\mu}\,\partial^\mu \delta F_{\alpha\beta}  \rangle \psi^\alpha \psi^\beta \nonumber\\&- \frac{i}{2}\langle \delta f \frac{\overleftarrow{\partial}}{\partial P^\mu} \delta F^{\mu\alpha} \rangle \psi_\alpha \chi\label{eq:scatteringterms}\,.
\end{align}
The equation for the fluctuations $\delta f$ is in general a complicated expression involving products of $\delta f$ and $\delta F^{\mu\nu}$
to higher orders. It serves as a generating functional for equations for higher moments of fluctuations, resulting in an infinite hierarchy 
of coupled equations, analogous to the BBGKY hierachy \cite{BBGKY,deGroot,Landau}. As noted previously, since  classical-statistical simulations are sufficient to describe single particle distributions for $\bar{f}\geq 1$, we can simplify the above to consider the dilute limit of $\delta f\ll \bar{f} \lessapprox 1$, where the collision term for the equation for fluctuations  is sub-dominant. Therefore keeping only linear terms in both  $\delta f$ and $\delta F^{\mu\nu}$, we obtain the evolution equation for $\delta f$ to be 
\begin{align}
\delta f\Big(\frac{\overleftarrow{\partial}}{\partial x^\mu}
\Big[\epsilon P^\mu +& \frac{i}{2}\psi^\mu \chi\Big]
\frac{\overleftarrow{\partial}}{\partial P^\mu}\Big[ \epsilon \bar{F}^{\mu\alpha}P_\alpha -\frac{i\epsilon}{2}\psi^\alpha \partial^\mu \bar{F}_{\alpha\beta}\psi^\beta \nonumber\\&+\frac{i}{2}\bar{F}^{\mu\alpha}\psi_\alpha \chi\Big]+\frac{\overleftarrow{\partial}}{\partial \psi^\mu}\Big[ \epsilon \bar{F}^{\mu\alpha}\psi_\alpha + \frac{P^\mu}{2}\chi\Big]+\frac{\overleftarrow{\partial}}{\partial \psi_5}\Big[ \frac{m}{2}\chi\Big] 
\Big)=K[\delta F]\,,
\label{Eq:lioudetails3} 
\end{align}  
where
\begin{align}
K[\delta F]\equiv -\bar{f}\Big(\frac{\overleftarrow{\partial}}{\partial P^\mu}\Big[ \epsilon \delta{F}^{\mu\alpha}P_\alpha -\frac{i\epsilon}{2}\psi^\alpha \partial^\mu \delta{F}_{\alpha\beta}\psi^\beta +\frac{i}{2}\delta{F}^{\mu\alpha}\psi_\alpha \chi\Big]
+\frac{\overleftarrow{\partial}}{\partial \psi^\mu}\Big[\epsilon \delta F^{\mu\alpha}\psi_\alpha\Big] \Big) \,.
\end{align}
The coupled set of equations \Eq{Eq:lioudetails2} and \Eq{Eq:lioudetails3} constitute the key ingredients of the chiral kinetic theory that we wish to construct. Much work remains to flesh out the structure of these equations and write them down in a fashion that is of practical use. This work is in progress. 

Missing from our preceding kinetic theory discussion is the contribution from the chiral anomaly, which was the principal motivation for this work. However we know how to include the anomaly in the SK framework since its contribution is contained in the fermion zero modes in the SK path integral. Finally, since the Liouville equation must be evaluated inside the world-line path integral, it is beneficial to split the distribution function $f$ into distributions for positive and negative chiralities. 
In the case of chiral fermions, the Hamiltonian in \Eq{eq:MassiveHamiltonian} is modified to include the Weyl constraints
\begin{align}
\frac{1}{2}(\gamma\cdot p)(1\pm \gamma^5)\Psi=0\,.\label{eq:WeylConstraint}
\end{align}
The conservation and non-conservation of chiral distribution functions can  then be understood
easily when the above discussion is generalized to the case of chiral Weyl fermions. Work in this direction is promising and will be reported on separately. 

\section{Acknowledgments}
The authors are supported by the U.S. Department of Energy,
Office of Science, Office of Nuclear Physics, under contract No. DE- SC0012704, and within the framework of the Beam Energy Scan Theory (BEST) Topical Collaboration.

\end{document}